The 8th term of an AP is 11 and the 15th term is 21. find the sum of the first 8 terms
1
Expert's answer
2019-11-05T10:39:57-0500

For the n-th term of AP we have


"a_n=a_1+(n-1)d"

Therefore


"a_8=a_1+7d=11"

"a_{15}=a_1+14d=21"

We obtain


"a_1=1,\quad d=10/7"

The sum of the first 8 terms


"S_8=\frac{a_1+a_8}{2}\times 8=\frac{1+11}{2}\times 8=48"